Find sin t, cos t, and tan t when the terminal side of an angle of t radians passes through the given point.
1) (-π, 2)

Answer by stanbon(75887)

Find sin t, cos t, and tan t when the terminal side of an angle of t radians passes through the given point. 1) (-π, 2)
x = -pi
y = 2
So, r = sqrt(pi^2 + 4)
------
sin = y/r = 2/sqrt(pi^2+4)
cos = x/r = -pi/sqrt(pi^2+4)
tan = y/x = -2/pi
---------------------
Cheers,
Stan H.
---------------------